OpenStudy (anonymous):

Slopes: Ok so my question is if you have (-1,-4) in your equation and your trying to see if the the coords work with each equation.... If one set is both negative and the other set has 1 of each 1 positive and 1 negative does that mean that both have to have addition signs?

OpenStudy (asnaseer):

I am not sure I understand what you are asking here. It might be better to give an example of a problem that shows what you mean.

OpenStudy (anonymous):

ok the coordinates to this equation that im doing are (-6,2) and (-1,-4) so the first one i found that worked was y + 4 = -6/5(x+1) would the next one have to be with addition signs or does it not matter?

OpenStudy (asnaseer):

If I understand your question correctly you are given two coordinates and are asked to find the equation of the line that passes through these two points - correct?

OpenStudy (anonymous):

yes Which of the following equations describe the line shown below?

OpenStudy (asnaseer):

in "the line shown below" - do they give you a graph of the line or just the two coordinates (-6,2) and (-1,-4)?

OpenStudy (anonymous):

they do yes. But trust me you dont want to see my drawing skills.

OpenStudy (asnaseer):

:)

OpenStudy (asnaseer):

OK - lets say you had two coordinates \((x1,y1)\) and \((x2,y2)\) and you were given a set of equations of straight lines. Then you can do this for each equation given: (1) plugin \(x=x_1\) into your equation and see if you get \(y=y_1\) as a result (2) plugin \(x=x_2\) into your equation and see if you get \(y=y_2\) as a result If both (1) and (2) result in the correct y values then that is the line that passes through both points

OpenStudy (asnaseer):

Now lets apply this to your equation:\[y+4=-\frac{6}{5}(x+1)\]

OpenStudy (asnaseer):

For \((-6,2)\) we have x=2 and y=-6. So lets replace x with 2 and see if we get y=-6:\[y+4=-\frac{6}{5}((2)+1)=-\frac{6}{5}(2+1)=-\frac{6}{5}\times3=-\frac{18}{5}\]\[\therefore y=-\frac{18}{5}-4=-\frac{38}{5}\ne-6\]therefore this equation does not represent a line that passes through the coordinates \((-6,2)\)

OpenStudy (anonymous):

ok.

OpenStudy (asnaseer):

For \((-1,-4)\) we have x=-4 and y=-1. So lets replace x with -4 and see if we get y=-1:\[y+4=-\frac{6}{5}((-4)+1)=-\frac{6}{5}(-4+1)=-\frac{6}{5}\times -3=\frac{18}{5}\]\[\therefore y=\frac{18}{5}-4=-\frac{2}{5}\ne-1\]therefore this equation does not represent a line that passes through the coordinates (−1,-4)

OpenStudy (anonymous):

i made a mistake it doesnt want the line passing through it wants the equations that describes the graph of the coordniates.

OpenStudy (asnaseer):

no problem - lets start again...

OpenStudy (asnaseer):

so you are asked to find the equation of a line that passes through: \((-6,2)\) and \((-1,-4)\) correct?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

In general, if you are given two coordinates, lets say \((x_1,y_1)\) and \((x_2,y_2)\) and are asked to find the equation of the line that passes through these two coordinates, then you can use this method:\[y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1)\]have you seen this equation before?

OpenStudy (anonymous):

not that particular one no

OpenStudy (asnaseer):

what have you been taught - I would rather try and teach you the same method that you use in class

OpenStudy (anonymous):

I take this online so I don't have a teacher.

OpenStudy (asnaseer):

ok - let me try and teach you the method above - if that is OK with you? :)

OpenStudy (anonymous):

Yes! I just want easy ways to figure this out so i can get the Slope unit knocked out and on to something new.

OpenStudy (asnaseer):

OK - a diagram will help, let me draw one...

OpenStudy (asnaseer):